OpenStudy (anonymous):

Solve the formula for f (g/f)+6=(5k)/f

OpenStudy (anonymous):

\[\large \frac{g}{f}+6=\frac{5k}{f}\]?

OpenStudy (anonymous):

if that is what you are given, multiply both sides of the equation by \(f\) and start with \[g+6f=5k\] and solve in two steps

OpenStudy (anonymous):

thanks!